# LLMBind: A Unified Modality-Task Integration Framework

### v3

@@ 3.1 LLMBind
3.2 LoRA Mixture-of-Experts
@@
$o=W_{0}x+\triangle Wx=W_{0}x+\sum_{i=1}^{N}G(x)_{i}E_{i}(x),$ (3)
$G(x)=Softmax(x\cdot W_{g}),$ (4)
@@
$o=W_{0}x+\gamma\sum_{i=1}^{N}G(x)_{i}E_{i}(x)=W_{0}x+\frac{\alpha}{r}\sum_{i=1}^{N}{w_{i}}B_{i}A_{i}x,$ (6)
@@
3.3 Training Objective
The training of our model involves three main objectives: text auto-regressive loss $\mathcal{L}_{reg}$, segmentation mask loss $\mathcal{L}_{mask}$ and MoE auxiliary loss $\mathcal{L}_{aux}$. These objectives are combined using weighted coefficients ($\lambda_{reg}$, $\lambda_{mask}$ and $\lambda_{aux}$) to compute the overall objective $\mathcal{L}$:
$\mathcal{L}=\lambda_{reg}\mathcal{L}_{reg}+\lambda_{mask}\mathcal{L}_{mask}+\lambda_{aux}\mathcal{L}_{aux}.$ (7)
@@ 3.3 Training Objective
$\mathcal{L}_{\text{reg}}(\theta)=\left\{\begin{array}[]{lr}-\sum_{i=1}^{N}log(p_{\theta}(x_{i}|X_{v},X_{t,<i})),&\text{if}X_{v}\subseteq X\\ \\ -\sum_{i=1}^{N}log(p_{\theta}(x_{i}|X_{t,<i})),&\text{if}X_{v}\nsubseteq X\\ \end{array}\right.$ (8)
@@
$\mathcal{L}_{mask}=\lambda_{bce}\mathcal{L}_{BCE}(\hat{{M}},{M})+\lambda_{dice}\mathcal{L}_{DICE}(\hat{{M}},{M}).$ (9)
@@
$\mathcal{L}_{aux}=\alpha\cdot N\cdot\sum_{i=1}^{N}f_{i}\cdot P_{i},$ (10)
@@
$f_{i}=\frac{1}{T}\sum_{x\in\mathcal{B}}\mathbbm{1}\{\text{argmax}\>p(x)=i\}.$ (11)
@@
3.4 Dataset
The training datasets consist of both publicly available datasets and our custom proprietary ones, as visualized in  Figure 5. These proprietary datasets are specifically constructed through interactive processes enabled by ChatGPT.
Figure 5: Different modality task datasets used in LLMBind, including VQA, Interactive Generation (video, audio, and image), Interactive Image Editing, Segmentation.
Interactive Image Generation and Editing Dataset. To empower the model in executing image generation and editing tasks, following Mini-DALLE3 (Zeqiang et al., 2023), we design a prompt for ChatGPT to construct 400k human-like interactive instruction data, tackling the problem of image generation and editing through textual description. As shown in  Figure 6, we first clarify the role and capability of ChatGPT within the prompt. We provide it with instruction text formats, tags, and examples for both generating and editing tasks. However, we notice that the generated samples closely resemble the examples in terms of text format, tone, and content. To address this, we further clarify the diversity of the target data in terms of the above multiple aspects in the prompt. Following a manual review of a portion of our dataset, we have confirmed that this approach is highly effective in ensuring significant diversity.
Table 2: Referring segmentation results on three referring segmentation datasets: refCOCO (Kazemzadeh et al., 2014), refCOCO+ (Kazemzadeh et al., 2014), and refCOCOg (Mao et al., 2016) with metric cIoU. The best results and second best results are indicated by boldface and underline, respectively.
@@ 3.4 Dataset
Table 3: Text-to-audio generation results on AudioCaps dataset (Kim et al., 2019).
Table 4: Text-to-video generation results on MSR-VTT dataset (Xu et al., 2016).
Table 5: Text-to-image generation results on COCO-captions dataset (Lin et al., 2014).
Method FD (normal-$\downarrow$) IS (normal-$\uparrow$)
DiffSound (Yang et al., 2023a) 47.68 4.01
AudioLDM-S (Liu et al., 2023b) 29.48 6.90
AudioLDM-L (Liu et al., 2023b) 23.31 8.13
NExT-GPT (Wu et al., 2023) 23.58 8.35
LLMBind (ours) 22.90 8.77
Method FID (normal-$\downarrow$) CLIPSIM (normal-$\uparrow$)
CogVideo (Hong et al., 2022a) 23.59 0.2631
Latent-VDM (Rombach et al., 2022a) 14.25 0.2756
Latent-Shift (An et al., 2023) 15.23 0.2773
CoDi (Tang et al., 2023) - 0.2890
NExT-GPT (Wu et al., 2023) 13.04 0.3085
LLMBind (ours) 11.09 0.2930
Method FID (normal-$\downarrow$)
@@
LLMBind (ours) 11.21
Table 4: Text-to-video generation results on MSR-VTT dataset (Xu et al., 2016).
Table 5: Text-to-image generation results on COCO-captions dataset (Lin et al., 2014).
Table 6: Reasoning segmentation results among LLMBind (ours) and previous related works. The phrase “1.5x7B” means that half of the layers in the 7B model are MoE layers.
Method gIoU cIoU
OVSeg (Liang et al., 2023) 28.5 18.6
GRES (Liu et al., 2023a) 22.4 19.9
X-Decoder (Zou et al., 2023a) 22.6 17.9
SEEM (Zou et al., 2023b) 25.5 21.2
LISA-7B (Lai et al., 2023) 52.9 54.0
LISA-Llama2-13B (Lai et al., 2023) 60.0 67.8
LISA-13B (Lai et al., 2023) 56.2 62.9
LLMBind-1.5x7B (ours) 62.4 66.9
Interactive Video and Audio Generation Dataset. To construct video and audio generation instruction data, we adapt the t2x(Wu et al., 2023) dataset and change its format to resemble that used for the image generation task. Some examples have been listed in  Appendix B.
Figure 6: Prompts for ChatGPT to construct interactive image generation and editing data.
@@
Visual Question Answering Datasets. We incorporates the LLaVA-Instruct-665k dataset (Liu et al., 2023d) for visual question-answering task. This dataset enables the training of models to comprehend, reason, and make decisions accurately to questions based on visual content.
4 Experiments
4.1 Experimental Configuration
@@
Training Specifications. During training, we employ 8 NVIDIA A100 GPUs (40G). The training scripts are based on the deepspeed (Rasley et al., 2020) engine. We use the AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of 0.0003 and weight decay of 0. The WarmupDecayLR learning rate scheduler is adopted, with a warmup iteration count of 100. The auto-regressive loss and segmentation loss have weights of 1.0 each. The BCE loss and DICE loss have weights of 2.0 and 0.5 respectively. We set the batch size per device to 2, and accumulate gradients over 10 steps, with a total of 15k steps.
Table 7: Ablation study about the rank of LoRA.
Table 8: Ablation study about the training steps.
Table 9: Ablation study about the number of MoE layers.
@@
Table 8: Ablation study about the training steps.
Table 9: Ablation study about the number of MoE layers.
Table 10: Ablation study about the number of experts and the number of Top-K activated experts.
@@
Evaluation Metrics. Our experiments utilize some key metrics for precise evaluation. Text-to-audio generation is assessed using Fréchet Distance (FD) for distribution similarity and Inception Score (IS) for quality and diversity. Text-to-video generation leverages Fréchet Inception Distance (FID) for content quality and CLIPSIM for textual alignment. In text-to-image generation, FID evaluates image distribution similarity. For segmentation, gIoU is defined by the average of all per-image Intersection-over-Unions (IoUs), while cIoU is defined by the cumulative intersection over the cumulative union.
4.2 Quantitative Results
Video & Audio & Image Generation. In the text-to-audio generation task, as shown in Table 5, LLMBind adopts CoDi (Tang et al., 2023) as the baseline generation model and achieves an FD score of 22.90 and an IS score of 8.77 on the AudioCaps (Kim et al., 2019) dataset, outperforming models like NeXT-GPT and AudioLDM. In the text-to-image generation task, as shown in Table 5, LLMBind employs Stable Diffusion (Rombach et al., 2022a) as the baseline generation model and attains an FID score of 11.21 on the COCO-caption (Lin et al., 2014) dataset, surpassing models like NeXT-GPT and CoDi. In the text-to-video generation task, as depicted in Table 5, LLMBind selects ModelScopeT2V (Wang et al., 2023) as the baseline generation model and obtains an FID of 11.09 and a CLIPSIM score of 0.2930 on the MSR-VTT (Xu et al., 2016) dataset, with a higher FID score and a slightly lower CLIPSIM score than NExT-GPT. In terms of generation tasks, since LLMBind does not perform lossy training on the generation models, it can be assumed that its generative capability largely depends on the chosen generation models. With sufficient computational resources, LLMBind can employ better-performing models to achieve superior generation results selectively.
Reasoning Segmentation. We evaluate the performance of LLMBind on the referring segmentation task. For a fair comparison, we use the LISA-7B model as our benchmark. As shown in Table 6, LLMBind achieves GIoU and CIoU scores of 62.4 and 66.9 on the ReasonSeg dataset (Lai et al., 2023), which exceeds LISA-7B with a large margin (8.8 and 12.9) and comparable with LISA-13B.
4.3 Ablation Analysis
We select the reasoning segmentation for ablation analysis, since it is conducive to quantitative analysis and comparison.
Rank of LoRA. As shown in Table 9, we conduct detailed experiments with different ranks of LoRA and found that our framework exhibits slight performance variation across various settings. Therefore, we chose a final rank of 8, for a balance between training cost and model performance.
Training Steps. Regarding the number of training iterations, we design experiments with four distinct settings: 5k, 6k, 7k, and 8k training steps. These variations were chosen to systematically evaluate the impact of training duration on model performance. As detailed in Table 9, our findings reveal a clear trend: longer iteration steps consistently yield improvements in both gIoU and cIoU metrics.
Number of MoE Layers. For the configuration of MoE layers, we conduct experiments with 16, 10, 8, and 0 layers, with the results presented in Table 9. The model’s performance generally improves with an increasing number of MoE layers. Due to limited model memory, we do not perform experiments with 32 MoE layers, but we believe that more MoE layers could further enhance the performance.
Number of All Experts and Top-K Activated Experts. In our MoE framework, the number of all experts and the number of Top-K activated experts are both crucial. In Table 10, we briefly explore how different numbers of experts and activated experts affect model performance. The findings align with our expectations: when the number of experts is fixed, more activated experts improve model performance; similarly, increasing the number of experts enhances performance when the number of activated experts is constant.
5 Conclusion and Future Direction
@@
Impact Statements
Broader Impacts
@@
 Appendix for LLMBind
Appendix A Interactive Architecture
A.1 Multi-Turn and Interactive
A.2 Interactive Multi-modal Generation
A.3 Interactive Image Editing and Refinement
A.4 Interactive Text Conversation
A.5 Joint Training Challenge
In our initial experiments, we observe that LLMBind rapidly excels in generating task-specific text prompts for various modality-tasks. This result aligns with our expectations due to LLMs’ established strengths in text comprehension and generation. However, a significant challenge arises when incorporating semantic-Embedding tokens in tasks like segmentation. LLMBind struggles to differentiate between image comprehension and segmentation tasks, often confusingly including the segmentation token <seg> in its responses, regardless of the simplicity of the text instructions. To address this, we increase the weight of the text generation loss, decrease the weight of the segmentation loss, and augment the proportion of VQA data in training and successfully tackle this issue. It can be seen that in the process of joint multi-task training, the loss of weight coefficients between multiple tasks and the proportion of training data for different tasks are issues that need to be explored in practice.
Appendix B Demo and Training Datasets
Figure 7: Examples of Image Generation and Editing.
Figure 8: Examples of Image Generation and Editing.
Figure 9: Examples of Image Segmentation.
Figure 10: Examples of Video Generation.
Figure 11: Examples of Audio Generation.

### v4

@@ 3.1. LLMBind
3.2. LoRA Mixture-of-Experts
@@
(3) $o=W_{0}x+\triangle Wx=W_{0}x+\sum_{i=1}^{N}G(x)_{i}E_{i}(x),$
(4) $G(x)=Softmax(x\cdot W_{g}),$
@@
(6) $o=W_{0}x+\gamma\sum_{i=1}^{N}G(x)_{i}E_{i}(x)=W_{0}x+\frac{\alpha}{r}\sum_{i=1}^{N}{w_{i}}B_{i}A_{i}x,$
@@
3.3. Training Objective
The training of our model involves three objectives: text auto-regressive loss $\mathcal{L}_{reg}$, segmentation mask loss $\mathcal{L}_{mask}$ and MoE auxiliary loss $\mathcal{L}_{aux}$. These objectives are combined using weighted coefficients ($\lambda_{reg}$, $\lambda_{mask}$ and $\lambda_{aux}$) to compute overall objective $\mathcal{L}$:
(7) $\mathcal{L}=\lambda_{reg}\mathcal{L}_{reg}+\lambda_{mask}\mathcal{L}_{mask}+\lambda_{aux}\mathcal{L}_{aux}.$
@@ 3.3. Training Objective
(8) $\mathcal{L}_{\text{reg}}(\theta)=\left\{\begin{array}[]{lr}-\sum_{i=1}^{N}log(p_{\theta}(x_{i}|X_{v},X_{t,<i})),&\text{if}\ X_{v}\subseteq X\\ \\ -\sum_{i=1}^{N}log(p_{\theta}(x_{i}|X_{t,<i})),&\text{if}\ X_{v}\nsubseteq X\\ \end{array}\right.$
@@
(9) $\mathcal{L}_{mask}=\lambda_{bce}\mathcal{L}_{BCE}(\hat{{M}},{M})+\lambda_{dice}\mathcal{L}_{DICE}(\hat{{M}},{M}),$
@@
(10) $\mathcal{L}_{aux}=\alpha\cdot N\cdot\sum_{i=1}^{N}f_{i}\cdot P_{i},$
@@
(11) $f_{i}=\frac{1}{T}\sum_{x\in\mathcal{B}}\mathbbm{1}\{\text{argmax}\>p(x)=i\}.$
@@
Table 2. Referring segmentation results on three referring segmentation datasets: refCOCO (Kazemzadeh et al., 2014), refCOCO+ (Kazemzadeh et al., 2014), and refCOCOg (Mao et al., 2016) with metric cIoU. The best results and second best results are indicated by boldface and underline, respectively.
@@
Table 3. Text-to-image generation results on COCO-captions dataset (Lin et al., 2014), evaluated with the Fréchet Inception Distance (FID) for image quality.
Method FID ($\downarrow$)
Specialized Models
@@
LLMBind-SD1.5 16.41
LLMBind-SD2 14.29
LLMBind-GALIP* 10.38
Figure 5. Different modality task datasets used in LLMBind, including VQA, Interactive Generation (video, audio, and image), Interactive Image Editing, and Segmentation.
Figure 6. Prompts for ChatGPT to construct interactive image generation and editing data.
3.4. Dataset
The training datasets consist of both publicly available datasets and our custom proprietary ones, as visualized in  Fig. 5. These proprietary datasets are specifically constructed through interactive processes enabled by ChatGPT.
Interactive Image Generation and Editing Dataset. To empower the model in executing image generation and editing tasks, following Mini-DALLE3 (Zeqiang et al., 2023), we design a prompt for ChatGPT to construct 400k human-like interactive instruction data, tackling the problem of image generation and editing through textual description. As shown in  Fig. 6, we first clarify the role and capability of ChatGPT within the prompt. We provide it with instruction text formats, tags, and examples for both generating and editing tasks. However, we notice that the generated samples closely resemble the examples in terms of text format, tone, and content. To address this, we further clarify the diversity of the target data in terms of the above multiple aspects in the prompt. Following a manual review of a portion of our dataset, we have confirmed that this approach is highly effective in ensuring significant diversity.
Interactive Video and Audio Generation Dataset. To construct video and audio generation instruction data, we adapt the t2x(Wu et al., 2023) dataset and change its format to resemble that used for the image generation task. Some examples are listed in Appendix C.
@@ 3.4. Dataset
Table 4. Text-to-audio generation results on AudioCaps dataset (Kim et al., 2019), evaluated with FD for audio similarity and IS for quality and diversity.
Method FD ($\downarrow$) IS ($\uparrow$)
Specialized Models
DiffSound (Yang et al., 2023b) 47.68 4.01
AudioLDM-S (Liu et al., 2023a) 29.48 6.90
AudioLDM-L (Liu et al., 2023a) 23.31 8.13
Multitask Models
NExT-GPT (Wu et al., 2023) 23.58 8.35
LLMBind-CoDi 22.90 8.77
LLMBind-Auffusion 21.99 10.57
(a) Text to video generation
(b) Image generation and edition
(c) Image segmentation
(d) Text to audio generation
Figure 7. Demonstration examples of the proposed LLMBind framework on different modality tasks.
@@
Visual Question Answering Datasets. We incorporates the LLaVA-Instruct-665k dataset (Liu et al., 2023d) for visual question-answering task. This dataset enables the training of models to comprehend, reason, and make decisions accurately based on visual content.
4. Experiments
4.1. Experimental Setting
@@
Training Specifications. During training, we employ 8 NVIDIA A100 GPUs (40G). The training scripts are based on the deepspeed (Rasley et al., 2020) engine. We use the AdamW (Loshchilov and Hutter, 2017) optimizer with a learning rate of 0.0003 and weight decay of 0. The WarmupDecayLR learning rate scheduler is adopted, with a warmup iteration count of 100. The auto-regressive loss and segmentation loss have weights of 1.0 each. The BCE loss and DICE loss have weights of 2.0 and 0.5 respectively. We set the batch size per device to 2, and accumulate gradients over 10 steps, with a total of 15k steps.
Evaluation Metrics. Our experiments utilize some key metrics for precise evaluation. Text-to-audio generation is assessed using Fréchet Distance (FD) for distribution similarity and Inception Score (IS) for quality and diversity. Text-to-video generation leverages Fréchet Inception Distance (FID) for content quality and CLIPSIM for textual alignment. In text-to-image generation, FID evaluates image distribution similarity. For segmentation, gIoU is defined by the average of all per-image Intersection-over-Unions (IoUs), while cIoU is defined by the cumulative intersection over the cumulative union. In addition, to better showcase the model’s generative capabilities during user interaction, we employed human evaluation to assess the quality of the modality data generated and its alignment with the given text instructions.
Table 5. Text-to-image generation results on COCO-captions dataset (Lin et al., 2014), evaluated with the Fréchet Inception Distance (FID) for image quality.
Method FID ($\downarrow$)
Specialized Models
CogVideo (Ding et al., 2021b) 27.10
GLIDE (Nichol et al., 2022) 12.24
CoDi (Tang et al., 2023) 11.26
Multitask Models
NExT-GPT (Wu et al., 2023) 11.28
LLMBind-SD1.5 16.41
LLMBind-SD2 14.29
LLMBind-GALIP* 10.38
4.2. Results Visualization
As shown in  Fig. 7, LLMBind can be executed on various modal tasks such as video, audio, image generation, image editing, and image segmentation, providing users with a more user-friendly interactive experience.
4.3. Quantitative Analysis
Image Generation. In the text-to-image generation task, as shown in Table 5, LLMBind utilizes various baseline models, achieving impressive FID scores on the COCO-caption (Lin et al., 2014) dataset. Specifically, with Stable Diffusion v1.5 and v2.0 (Rombach et al., 2022c) as baselines, LLMBind attains FID scores of 16.41 and 14.29, respectively. Notably, using GALIP (Tao et al., 2023) as the baseline model, LLMBind secures an FID score of 10.38, outperforming models such as NeXT-GPT (Wu et al., 2023) and CoDi (Tang et al., 2023).
Audio Generation. In the domain of text-to-audio generation, LLMBind has been meticulously evaluated and compared with existing models. As depicted in  Table 4, LLMBind utilizes the CoDi (Tang et al., 2023) model as a baseline for audio generation, attaining an FD score of 22.90 and an IS of 8.77 on the AudioCaps (Kim et al., 2019) dataset. Furthermore, when employing Auffusion (Tang et al., 2023) as the foundational model for generation, LLMBind secures an FD score of 21.99 and an IS score of 10.57 on the same dataset, demonstrating a discernible advantage over models such as NeXT-GPT (Wu et al., 2023) and AudioLDM (Liu et al., 2023a).
Table 6. Text-to-video generation results on MSR-VTT dataset (Xu et al., 2016), evaluated with FID for video quality and CLIPSIM for video coherence.
Method FID ($\downarrow$) CLIPSIM ($\uparrow$)
Specialized Models
CogVideo (Hong et al., 2022b) 23.59 0.2631
Latent-VDM (Rombach et al., 2022a) 14.25 0.2756
Latent-Shift (An et al., 2023) 15.23 0.2773
Make-A-Video (Singer et al., 2022a) 13.17 0.3049
Multitask Models
NExT-GPT (Wu et al., 2023) 13.04 0.3085
LLMBind-Show-1 13.08 0.3072
LLMBind-ModelScopeT2V 11.09 0.2930
LLMBind-HiGen 8.60 0.2947
Table 7. Human evaluation statistics regarding the preference rate for Quality Score (Q.S.) and Alignment Score (A.S.). The former metric assesses the quality of generated content, while the latter evaluates the alignment between the generated data and textual instructions. Both metrics are rated on a scale of 1-4, with the final score being the weighted average.
Method Q.S. Final A.S. Final
1 2 3 4 score 1 2 3 4 Score
Audio Generation
Direct input 7% 35% 34% 24% 2.75 18% 39% 23% 20% 2.45
Ours 8% 28% 35% 29% 2.85 6% 24% 27% 43% 3.07
Image Generation
Direct input 5% 39% 43% 13% 2.64 21% 29% 45% 5% 2.34
Ours 5% 27% 41% 27% 2.90 12% 19% 51% 18% 2.75
Video Generation
Direct input 29% 41% 23% 7% 2.08 23% 42% 21% 14% 2.26
Ours 24% 38% 27% 11% 2.25 19% 40% 24% 17% 2.39
Video Generation. For the text-to-video generation task, LLMBind’s performance on the MSR-VTT (Xu et al., 2016) dataset is shown in Table 6. Using ModelScopeT2V (Wang et al., 2023) as the base model, LLMBind achieves an FID of 11.09 and a CLIPSIM score of 0.2930. With Show-1 (Zhang et al., 2023) as the base model, it obtains an FID of 13.08 and a CLIPSIM score of 0.3072. When employing HiGen (Qing et al., 2023) as the base model, LLMBind reaches an FID of 8.60 and a CLIPSIM score of 0.2947, showing comparable performance to NExT-GPT (Wu et al., 2023). As LLMBind avoids lossy training on generation models, its generative power primarily relies on the selected base models.
4.4. Human Evaluation
As shown in  Fig. 8, to assess the practical performance of LLMBind in interactive Image, Video, and Audio generation tasks (IVA), we conducted a comprehensive user study. This study is designed to measure the capability of models to generate IVA data in response to interactive text instructions, emulating natural human-like communication. The study focuses on assessing two critical metrics: Quality Score (Q.S.) and Alignment Score (A.S.). The Q.S. assesses the quality of the generated IVA content, while A.S. evaluates the alignment between the generated data and the textual instructions provided. Participants interact with the model by issuing text instructions and then rate each generated sample on a scale of 1 to 4 for both metrics, with 1 representing poor performance and 4 indicating excellent performance. The aggregated results from a panel of 30 participants, as illustrated in  Table 7, clearly indicate that LLMBind excels in both the quality of modal data generation and its ability to align with text-based instructions closely. The final scores, calculated as the weighted average of the ratings for Q.S. and A.S., demonstrate LLMBind’s proficiency in understanding and executing complex, multimodal tasks as directed by human users.
Figure 8. Human evaluation. The evaluation indicators for interactive generation include Q.S. and A.S.
Table 8. Ablation study about the rank of LoRA.
@@ 4.4. Human Evaluation
Table 9. Ablation study about the training steps.
@@
Table 10. Ablation study about the number of MoE layers.
@@
Table 11. Ablation study about the number of experts and the number of Top-K activated experts.
@@
4.5. Ablation Analysis
We apply the ablation analysis on the reasoning segmentation task, since it is conducive to quantitative analysis and comparison. The detailed tables are listed in the supplementary.
Rank of LoRA. As shown in  Table 8, our framework exhibits slight performance variation across different settings.
Training Steps. As shown in  Table 9, we reveal a clear trend that longer iteration steps consistently yield improvements in both gIoU and cIoU metrics.
Number of MoE Layers. As shown in  Table 10, model performance generally improves with the increase in the number of MoE layers.
Number of All Experts and Top-K Activated Experts. As shown in  Table 11, when the number of experts is fixed, more activated experts improve model performance. Conversely, increasing the number of experts enhances performance when the number of activated experts is constant.
5. Conclusion and Future Direction
@@
 Appendix for LLMBind
Appendix A Interactive Architecture
A.1. Multi-Turn and Interactive
A.2. Interactive Multi-modal Generation
A.3. Interactive Image Editing and Refinement
A.4. Interactive Text Conversation
Appendix B Impact Statements
B.1. Broader Impacts
@@ B.3. Licenses
Appendix C Demo and Training Datasets
Figure 9. Example of Image Generation and Editing (1).
Figure 10. Example of Image Generation and Editing (2).
Figure 11. Example of Image Generation and Editing (3).
Figure 12. Example of Image Generation and Editing (4).
Figure 13. Example of Image Segmentation (1).
Figure 14. Example of Image Segmentation (2).
Figure 15. Example of Image Segmentation (3).
Figure 16. Examples of Video Generation.
Figure 17. Examples of Audio Generation.